The locus of the complex number arg(z* + i√3) - π/4 is equal to?


1
Expert's answer
2021-10-27T14:02:45-0400

Let "z=x+iy." Hence "x+i(y+\sqrt{3})=\lambda."

If "\arg(z+i\sqrt{3})=\pi/4," then


"x=y+\sqrt{3}=>y=x-\sqrt{3}"

The locus of the compex number "z" is equal to


"y=x-\sqrt{3}"


If "\arg(z^*+i\sqrt{3})=\pi/4," then "x+i(-y+\sqrt{3})=\lambda."


"x=-y+\sqrt{3}=>y=-x+\sqrt{3}"

The locus of the compex number "z" is equal to


"y=-x+\sqrt{3}"